The sides of an equilateral triangle are decreasing at a rate of 3 in/hr. How fast is the area enclosed by the triangle decreasing when the sides are 2 feet long?
1
Expert's answer
2020-07-27T19:03:55-0400

1 foot =12 inches, thus 2 feet =24 inches.

An equilateral triangle has all its side equal i.e s.

h= "\frac{✓3} {2}"s

A="\frac {1}{2}×s×\frac{✓3} {2}"

="\frac{1} {4}✓3"s2

If "\frac{ds} {dt}" =-3in/hr,when s=24, "\frac{dA}{dt}" =?

Differentiating with respect to t;

"\frac{dA} {dt}" "= \frac{✓3} {4}"×2s "\frac{ds} {dt}"

Hence s=24 when "\frac{dA} {dt}" ="\frac{✓3} {2}" (24)(-3)=-36✓3 in/hr


=-36✓3 in/hr